Let T be the linear operator on R


2 defined by

T(x, y) = (−y, x)

i. What is the matrix of T in the standard ordered basis for R2 ?

ii. What is the matrix of T in the ordered basis B = {α1, α2}, where α1 = (1, 2) and α2 = (1, −1)?

iii. Prove that for every real number c the operator (T − cI) is invertible.


1
Expert's answer
2022-05-06T10:47:08-0400

i. The operator "T(x,y)" can be written in the form: "\left(\begin{array}{ll}-y\\ x\end{array}\right)=\left(\begin{array}{ll}0&-1\\1&0\end{array}\right)\left(\begin{array}{ll}x\\ y\end{array}\right)". "\left(\begin{array}{ll}0&-1\\1&0\end{array}\right)" is the matrix of the operator in the standard ordered basis. The matrix can be easily found by considering equality: "\left(\begin{array}{ll}-y\\ x\end{array}\right)=\left(\begin{array}{ll}a_{11}&a_{12}\\a_{21}&a_{22}\end{array}\right)\left(\begin{array}{ll}x\\ y\end{array}\right)". From the latter we get: "-y=a_{11}x+a_{12}y" and "x=a_{21}x+a_{22}y". We immediately receive: "a_{11}=0,a_{12}=-1" and "a_{21}=1,a_{22}=0."

ii. In order to find the matrix in the ordered basis "B=\{ \alpha_1,\alpha_2\}", we consider the vector "\left(\begin{array}{ll}x\\ y\end{array}\right)" in the basis "B". We receive: "\left(\begin{array}{ll}x\\ y\end{array}\right)=\left(\begin{array}{ll}1\\ 2\end{array}\right)t_1+\left(\begin{array}{ll}1\\ -1\end{array}\right)t_2". We receive: "x=t_1+t_2", "y=2t_1-t_2". Thus, "t_1=\frac{x+y}{3}", "t_2=\frac{2x-y}3". In the same way we consider the vector "\left(\begin{array}{ll}-y\\ x\end{array}\right)=\left(\begin{array}{ll}1\\ 2\end{array}\right)s_1+\left(\begin{array}{ll}1\\ -1\end{array}\right)s_2". We receive: "s_1=\frac{-y+x}{3}", "s_2=\frac{-2y-x}{3}". The aim is to find the matrix satisfying: "\left(\begin{array}{ll}\frac{-y+x}{3}\\ \frac{-2y-x}{3}\end{array}\right)=\left(\begin{array}{ll}b_{11}&b_{12}\\b_{21}&b_{22}\end{array}\right)\left(\begin{array}{ll}\frac{x+y}{3}\\ \frac{2x-y}3\end{array}\right)". We receive: "\frac{-y+x}{3}=b_{11}\frac{x+y}{3}+b_{12}\frac{2x-y}3", "\frac{-2y-x}{3}=b_{21}\frac{x+y}{3}+b_{22}\frac{2x-y}3". We set "y=-x" in the first equation and receive: "b_{12}=\frac23". Setting "y=2x" yields: "b_{11}=-\frac{1}3". We set "y=-x" in the second equation and get: "b_{22}=\frac13". Setting "y=2x" provides: "b_{21}=-\frac53". Thus, matrix "T" in the ordered bases "B" has the form: "\left(\begin{array}{ll}-\frac13&\frac23\\-\frac53&\frac13\end{array}\right)".

iii. Matrix "T-cI" has the form: "T-cI=\left(\begin{array}{ll}-c&-1\\1&-c\end{array}\right)". "det(T-cI)=c^2+1>0". Thus, the matrix is invertible for all real "c". It means, that the operator is invertible.

Answer: i. "\left(\begin{array}{ll}0&-1\\1&0\end{array}\right)"; ii. "\left(\begin{array}{ll}-\frac13&\frac23\\-\frac53&\frac13\end{array}\right)"; iii. "det(T-cI)=c^2+1>0". Therefore, the operator is invertible.